if two angle of a triangle are equal and third angle is 50 degree find the two angles

Grade:7

4 Answers

Arun
4 years ago
 
Dear student
 
Let two equal angles are x and x
hence sum of all angles. = 180
x + x + 50 = 180
 
2x = 130
x = 65
Yashraj
3 years ago
let unknown angle is x
so 2x +50=180
sumof all anglesin triangle is 180
x=130/2
=65
…..................
Janarth
3 years ago
Let 2 of the equal angles be a and a
Sum of all angles will always add up to = 180

so, a+a+50 will be 180
2a = 180-50 = 130

2a=180

so,a will be 130/2  which is 65.
 
hope it helps.
Shubhanshu agrawL
3 years ago
Let the unknown angle be x, 
We know that sum of three angles of triangle is 180°
We have two angles is x and the last one is 50°
So, 
x+x+50°=180°
x= 65°